### v2

# Cost and accuracy of long-term memory in Distributed Multi-Agent Systems based on Large Language Models

Abstract
Distributed multi-agent systems (DMAS) based on large language models (LLMs) enable collaborative intelligence while preserving data privacy. However, systematic evaluations of long-term memory under network constraints are limited. This study introduces a flexible testbed to compare mem0, a vector-based memory framework, and Graphiti, a graph-based knowledge graph, using the LoCoMo long-context benchmark. Experiments were conducted under unconstrained and constrained network conditions, measuring computational, financial, and accuracy metrics. Results indicate mem0 significantly outperforms Graphiti in efficiency, featuring faster loading times, lower resource consumption, and minimal network overhead. Crucially, accuracy differences were not statistically significant. Applying a statistical Pareto efficiency framework, mem0 is identified as the optimal choice, balancing cost and accuracy in DMAS.
Keywords: distributed multi-agent systems; large language models; long-term memory; mem0; Graphiti; knowledge graphs; vector databases; cost-accuracy trade-off; LLM benchmarking
Open-source research: the source code of the testbed, experimental results as well as notebooks for analyzing and evaluating the results are available at: https://github.com/wolffbe/dmas-memory.
List of acronyms
Acronyms
AES
accuracy-efficiency score
AI
artificial intelligence
API
application programming interface
AWS
Amazon Web Services
CI
confidence interval
CpSQ
cost per successful query
CPU
central processing unit
DMAS
distributed multi-agent system
IDK
i don’t know
JSON
JavaScript object notation
LLM
large language model
LTS
long-term support
MAS
multi-agent system
Mbits
megabits
MBytes/s
megabyte per second
Q&A
questions and answers
RAG
retrieval-augmented generation
RAM
random access memory
SLM
small language model
TCO
total cost of ownership
TCP
transmission control protocol
USD
United States Dollar
vCPU
virtual central processing unit
WSL
Windows Subsystem for Linux
1 Introduction
Artificial intelligence (AI) agents based on large language models (LLMs) are capable of methodically translating goals into actions, accessing tools interfacing third-party services, and learning from past experiences autonomously retrieved from memory [87].
By connecting multiple LLM-based agents, multi-agent systems (MASs) achieve a higher form of collective intelligence that has been applied in various fields and industries [22, 59, 85, 62, 32, 30, 17], e.g. poetry [84], construction [34] and finance [71].
Distributed multi-agent systems (DMASs) emerged from a need to mitigate privacy risks as centralized parties in MASs process sensitive user information [75]. The vision is a global network of DMASs that collaborate using local private memory to complement knowledge-intensive human labor, research and innovation [3].
Long-term memory benchmarks such as LoCoMo exist to evaluate the contextual learning ability of LLMs [38] and MASs [68]. Various memory providers have already proven to enhance contextual learning using such benchmarks [68, 49, 39, 12]. Recent research has also been focused on identifying best practices and agent design patterns to increase the cost-effectiveness of MASs [73].
However, systematically evaluating the contextual learning capabilities of novel memory frameworks combined with newly researched best practices and design patterns for MASs remains challenging. Memory frameworks are often assessed solely on accuracy and token usage without integration into actual MASs, overlooking additional computational costs and latency. New design patterns are typically evaluated in domain-specific implementations that lack repeatable and comparable performance tests across different MASs and benchmarks. Finally, no research exists so far that systematically examines and enhances long-term memory in DMASs, particularly across different network scenarios.
1.1 Theoretical framework and literature study
A review of 67 papers addressing LLM-based agents, MASs and DMASs revealed five common fields of research across all three agent types, namely technology, optimization, safety and security, networking and memory. No existing literature on safety and security for LLM-based agents and memory for DMASs was identified throughout the review. The following chapter highlights the theoretical framework and literature that provide the foundation to this, presumably, first research of memory for DMASs.
LLM-based agents
The four components of LLM-based agents are perception, the ability to interact with an environment and receive feedback, planning, the ability to divide tasks into smaller sub-tasks, memory, the ability to persist an understanding of sources and operations over time, and action, the ability to take necessary steps to reach a given goal [14]. A second approach similarly divides an LLM-based agent’s structure into profile, perception, mutual interaction, and evolution [33].
Optimizing agent effectiveness ranges from parameter-driven methods such as fine-tuning to parameter-free approaches such as prompt engineering and external knowledge retrieval [15]. A third parameter-free approach is chain-of-thought reasoning enabling an LLM to structure complex thought processes [21]. Additionally, the cost of an agent is optimized by decomposing multi-LLM selection challenges into a relaxed form [13], extracting and caching plan templates from successful task completions [83] or adjusting the complexity of a given task [64].
Use cases for agents in networking are network slicing [60], viewport prediction, adaptive bitrate streaming, cluster job scheduling [70], optimizing networks in Open Radio Access Network environments [52], mitigating data exfiltration scenarios [51], and managing software-defined networks [5]. Another use case is the distributed inference of LLMs at the edge of a network by dividing the model into partitions, which are then assigned to different devices [37], for example, mobile phones [42]. The placement and migration of such distributed LLMs have also been managed using ant colony algorithms [67].
Literature on agent memory consists of frameworks and benchmarks. Memory frameworks are categorized into vector storages such as VIMBank [31] and mem0 [12] as well as graph architectures such as Zep [49]. HiAgent is an example of a third group of frameworks that uses hierarchical working memory to approach tasks that require large amounts of memory [24]. A fourth approach is retrieval-augmented generation (RAG) that extends vector storages by combining dense vectors with parametric memory based on a pre-trained sequence-to-sequence model [4].
Memory benchmarks evaluate the performance of memory frameworks. For example, LoCoMo and LongMemEval provide dialogues to assess temporal reasoning [38, 69], LongBench v2 tests comprehension and reasoning over contexts ranging from eight thousand to two million words [7], and StoryBench uses interactive fiction with branching storylines to examine hierarchical decision-making [63].
Multi-agent systems
MASs are categorized based on actors, types, communication structures, collaboration strategies, and coordination mechanisms [61]. Structures range from centralized to fully decentralized as well as static to dynamic implementations [74]. Similarly to agent memory benchmarks, MAS structure benchmarks such as AgentsNet exist to assess their performance [20].
DyLAN offers a two-stage approach to actor optimization that consists of an agent selection algorithm prior to establishing a dynamic collaboration strategy [36]. A second optimization approach involves categorizing agents into planners, reasoners and verifiers followed by identifying redundant reasoning results [65]. Additionally, profiling, efficient retrieval of past observations, LLM cascades and ask-the-expert calls have been implemented to optimize cost [18]. A third approach involves graph networks to find the optimal communication structure [76], and optimize asynchronous and parallel operations [81].
Safety and security is named as a key research direction in the field of MASs [58]. Failures in MASs have led to specification issues, for example, disobeying task or role specifications, inter-agent misalignment such as failing to ask for clarification, withholding information or ignoring other agent’s inputs, and task verification, e.g. premature termination or incomplete, incorrect or no verification [9]. Moreover, adversarial debates, voting mechanisms and error logs mitigate the risk of agents hallucinating [73]. Graph networks have also been employed in this context to detect and defend malicious agents and implement structural interventions [66]. In general, chain and cycle structures offer a high degree of protection against the spread of malicious information [77].
Use cases of MASs managing other networks include anomaly detection [46], simulations [50], mediating multi-party business and network goals [10] and orchestrating one or more centralized [19] or decentralized [29] networks [47]. Orchestration approaches range from deploying mixture of experts architectures [56] and multi-agent reinforcement learning [82] handling data retrieval, collaborative planning as well as evaluation and reflection [28].
Short-term memory in MASs is defined as only existent for the duration of the ongoing interaction while long-term memory consists of historical queries and responses [23]. Similarly to LLM-based agents, hierarchical MAS memory such as G-Memory [80] and graph-based frameworks such as AGENTiGraph exist [88]. In the field of MASs long-term memory, MIRIX combines core, episodic, semantic, procedural, resource memory and knowledge vault memory with a MAS that manages knowledge updates and retrieval [68]. Alternatively, Intrinsic Memory Agents construct agent-specific heterogeneous structured memories that evolve with agent outputs [78].
Distributed multi-agent systems
An initial implementation of a DMAS consists of a cloud-based LLM-based agent as a central planner in combination with device-specific agents and generative agents to optimize task execution efficiency and accuracy [72]. However, different networking scenarios were not further considered.
A first approach to optimizing multi-agent systems in edge scenarios is Context-Adaptive Sparse Key-Value, which identifies the most relevant context tokens and prunes less critical key-value pairs based on usage frequency and recency [41].
In regard to DMAS safety and security, the four most critical trustworthiness challenges include free riding, susceptibility to malicious attacks, communication inefficiencies, and system instability [86]. Consequently, security measures on an agent or intra-LLM level and system- or inter-LLM level are required. Intra-LLM security consists of strong LLM identity, authentication and authorization, context-aware access control as well as stateless and ephemeral LLM management. Inter-LLM security covers proactive maintenance, blockchain and distributed management, micro-segmentation and isolation as well as intelligent system monitoring and failure management [35].
Extending existing network management use cases in the field of MAS [48], AgentRAN proposes the AI-RAN Factory that generates new agents, enabling a distributed network to continuously develop and grow its own network of intelligent agents [16].
1.2 Research questions and hypothesis
Based on the theoretical framework and literature study, two research gaps were identified. First of all, and this issue prevails across all three agent system types, no testbed exists to comparatively evaluate the computational and financial cost of long-term memory across implementations in addition to token usage and benchmarks. Secondly, no DMAS has been built and tested taking different network scenarios into account. Consequently, an initial prototype of a MAS and DMAS long-term memory testbed is proposed, and applied to answer the following research questions:
  • Which memory framework provides the best balance between knowledge retention as well as computational and financial cost in MASs?
  • How do knowledge retention as well as computational and financial cost vary when MASs operate in a hybrid cloud–edge environment?
2 Method
An empirical approach was employed [8]. Two memory frameworks were selected for comparative evaluation: mem0, a vector-based system with LLM compression [12] that recently secured \$24 million United States Dollar (USD) in funding [55], and Graphiti, a graph database based on Neo4j, the latter having raised \$325 million USD in Series F investment, representing the largest funding round in database history [43].
2.1 Empirical approach
Experiments were conducted employing a factorial design with two independent variables: the memory framework and the network profile. The unconstrained experiment added no artificial latency or bandwidth restrictions between agents, allowing evaluation under ideal network conditions. The constrained experiment introduced realistic limitations, including 200 milliseconds of latency, a bandwidth limit of eight megabits (Mbits) or one megabyte per second (MBytes/s), and 50 milliseconds of jitter, simulating a typical datacenter-to-edge connection.
The LoCoMo long-context memory benchmark, previously used in the initial evaluation of mem0 and MIRIX [68, 12], served as the source of context. Each experiment consisted of two phases, the loading and the questions and answers (Q&A) phase. During the loading phase, all 19 sessions contained in the first of ten conversations from LoCoMo were added to mem0 and Graphiti. For reference, the JavaScript object notation (JSON) structure of a session is presented in appendix A.1. During the Q&A phase, the MAS answered 199 questions provided in LoCoMo about all 19 sessions. By adding all sessions from the first conversation, and asking all questions related to the first conversation, a robust sample size with statistical relevance was ensured.
While loading the conversation into memory and answering questions, execution time, central processing unit (CPU) and random access memory (RAM) usage on edge and cloud, disk and network utilization, OpenAI token consumption and associated cost, application programming interface (API) request latency, the actual and received answers, as well as their similarity, were recorded as ten dependent variables.
Two types of similarity were evaluated: string similarity, computed using SentenceTransformer to assess textual alignment between expected and received answers [26], and semantic similarity, measured using the transformer-based embedding model all-MiniLM-L6-v2 [25]. The final similarity score for each response was calculated as the average of these two measures.
2.2 Testbed architecture
The testbed runs on Ubuntu inside a Windows Subsystem for Linux (WSL) that is installed on a physical machine containing an Intel Core processor and running Windows as a native operating system. The MAS is composed of three Docker containers, each running Python. During the loading phase, the locomo service loads the LoCoMo benchmark dataset from GitHub [57], and exposes an API that enables the coordinator to request individual conversations and sessions, and load them into memory. For that, the coordinator decomposes conversations and sessions into individual turns and sends them to the memory agent via its /remember endpoint.
The memory container operates in its own memory-internal subnet that also hosts the qdrant Docker container with the volume qdrant-data [40], and the neo4j Docker container with the volumes neo4j-logs and neo4j-data [79]. Furthermore, mem0, based on qdrant and Graphiti, based on neo4j are both defined as services in the memory agent. Based on the configuration in the repository’s .env file, the memory agent persists memories using either the mem0 or Graphiti backend. All other testbed components run in the DMAS-network.
Questions about each conversation are issued through the coordinator’s /ask endpoint. As a result, the coordinator forwards each question to the memory agent to retrieve related memories, and sends the question together with the retrieved context to the responder. The responder then returns its answer to the coordinator, which returns the final response to the APIs analysis Jupyter notebook that collects all questions, answers and metrics.
The coordinator relies entirely on tools that are called by the small language model (SLM) qwen2.5:3b-instruct-q4_K_M with three billion parameters that is optimized for tool calling [45]. On the other hand, the responder uses the LLM gpt-4o-mini with approximately eight billion parameters. The system and user prompts that are invoked during this process are shown in appendix A.2, and an example of retrieved memories sent to the responder is presented in appendix A.3.
The SLM is hosted in ollama [44], and stored in the ollama-data volume. Both the memory agent and the responder integrate the official OpenAI Platform. All three agents route their LLM requests through the ollama-proxy and openai-proxy, so token usage is tracked in prometheus, which stores all metrics collected over time in the prometheus-data volume.
CPU, RAM, disk and network metrics are collected using telegraf [27]. The docker-proxy based on socat exposes the Docker Unix socket over a transmission control protocol (TCP) port allowing telegraf to access the Docker daemon [1].
All traffic between the coordinator, memory, and the responder is routed through toxiproxy, which applies artificial latency, bandwidth and jitter [54] in case of running the constrained experiment.
Connecting all components, a Makefile automates the testbed setup, and executes the analysis Jupyter notebook that loads conversations, queries the MAS, and stores all results in CSV files, with one CSV generated for each combination of memory framework and experiment. A detailed overview of testbed components and their versions is located in appendix A.4.
Figure 1: testbed system architecture
3 Results and analysis
During the unconstrained experiment, Graphiti required 3.96% more OpenAI tokens than mem0. During the constrained experiment, Graphiti cost 32% more tokens. Table 1 presents the financial cost including total tokens and the corresponding cost in USD.
experiment phase memory tokens ($\times 10^{3}$) cost (USD)
unconstrained loading Graphiti 4,948.5 \$0.0027
mem0 1,017.4 \$0.1319
Q&A Graphiti 161.5 \$0.4287
mem0 113.2 \$0.2822
constrained loading Graphiti 4,956.7 \$0.0030
mem0 1,014.2 \$0.1319
Q&A Graphiti 112.5 \$0.2775
mem0 113.4 \$0.2806
Table 1: financial cost per experiment, phase, and memory
Each component in the system was labelled as cloud or edge to understand which network component consumed more computational resources, especially in consideration of answering to the second research question. The coordinator, ollama-proxy, ollama and ollama-data were labelled as cloud components. The responder, openai-proxy, memory, qdrant, qdrant-data, neo4j-logs and neo4j-data were marked as edge components.
During the unconstrained experiment, Graphiti used 104.3% more CPU in the cloud and 232.5% more at the edge, 38.7% more RAM at the cloud and 8% more at the edge, 517.4% more disk in the cloud and 88.9% less at the edge, and 6,283.5% more network bandwidth in the cloud and 3,966.7% more at the edge than mem0.
During the constrained experiment, Graphiti used 77.2% more CPU in the cloud and 80.7% more at the edge, 54.8% more RAM in the cloud and 24% more at the edge, 430.7% more disk in the cloud and 89.7% less at the edge, and 5,268.7% more network bandwidth in the cloud and 4,295% more at the edge than mem0. Table 2 details the specific usage of computational resources across both network environments, broken down by memory framework, experiment phase, and cloud versus edge consumption.
cloud edge
CPU RAM disk network CPU RAM disk network
experiment phase memory (minutes) (MB) (MB) (MB) (minutes) (MB) (MB) (MB)
unconstrained loading Graphiti 17.1 1,115.3 76.0 1,332.3 3.2 145.3 0.0 71.1
mem0 2.4 766.9 11.8 22.1 0.4 94.5 0.0 1.7
Q&A Graphiti 2.1 1,221.2 5.5 97.4 10.1 2,617.1 125.0 2.1
mem0 7.0 918.3 1.4 0.3 3.6 2,463.8 1,122.9 0.1
constrained loading Graphiti 6.2 1,726.5 84.4 1,335.9 1.2 196.8 0.1 87.7
mem0 3.3 826.7 10.5 24.1 0.6 96.3 0.1 1.9
Q&A Graphiti 11.7 914.7 2.1 0.7 9.1 3,229.5 46.3 0.2
mem0 6.8 879.8 5.8 0.8 5.1 2,667.1 449.1 0.1
Table 2: computational cost per experiment, phase, and memory
Based on Table 3 below displaying the total time for each phase of both experiments and memory frameworks, mem0 was 86.5% faster during the loading phase and 14.7% faster during the Q&A phase than Graphiti. During the constrained experiment, mem0 was 82.8% faster during the loading phase and 14.5% faster during the Q&A phase than Graphiti.
experiment phase memory duration (minutes)
unconstrained loading Graphiti 180.32
mem0 24.33
Q&A Graphiti 53.77
mem0 45.90
constrained loading Graphiti 173.95
mem0 30.01
Q&A Graphiti 56.03
mem0 47.90
Table 3: temporal cost per experiment, phase, and memory
Regarding knowledge retention and efficiency, the responder’s prompt was adjusted to answer with i don’t know (IDK) rather than being indecisive or inventing a response. As a result, Graphiti was 3.6% more accurate than mem0 in the unconstrained experiment. During the constrained experiment, Graphiti was 2% more accurate.
Table 4 details the absolute numbers and rates for correct, wrong, and indecisive answers. Furthermore, Figure 2 displayed on the next page visualizes these distributions, highlighting that while Graphiti offers slightly improved accuracy, both systems are dominated by conservative IDK responses.
results rates (%)
experiment memory questions correct wrong IDK accuracy IDK answers
unconstrained Graphiti 199 22 60 117 11.1% 58.8% 41.2%
mem0 199 15 53 131 7.5% 65.8% 34.2%
constrained Graphiti 199 16 50 133 8.0% 66.8% 33.2%
mem0 199 12 51 136 6.0% 68.3% 31.7%
Table 4: accuracy per experiment and memory
Figure 2: response distribution (Correct, Wrong, IDK) for mem0 and Graphiti under unconstrained and constrained network conditions.
To determine whether the observed performance differences represent a true systemic advantage or merely sampling artifacts, two statistical methods were employed: 95% Wilson confidence intervals (CIs) for individual accuracy rates and two-proportion z-tests for pairwise comparisons.
Table 5 presents the accuracy rates with their corresponding confidence intervals. Notably, the intervals for Graphiti and mem0 overlap substantially in both profiles, providing an initial indication that their performance distributions are not distinct.
memory experiment accuracy 95% CI
Graphiti unconstrained 0.111 [0.074, 0.162]
mem0 unconstrained 0.075 [0.046, 0.121]
Graphiti constrained 0.080 [0.050, 0.127]
mem0 constrained 0.060 [0.035, 0.102]
Table 5: accuracy rates and 95% Wilson CIs
This lack of distinction is confirmed by the z-test results detailed in table 6. As shown, all $p$-values exceed the significance threshold of $\alpha=0.05$, meaning the null hypothesis regarding accuracy differences cannot be rejected.
comparison z-stat p-value
Graphiti vs mem0 (unconstrained) 1.208 0.2269
Graphiti vs mem0 (constrained) 0.784 0.4330
Graphiti: unconstrained vs constrained 1.023 0.3061
mem0: unconstrained vs constrained 0.598 0.5498
Table 6: two-proportion z-test results with no significant differences found at $\alpha=0.05$
While empirical data suggests a positive trend in favor of Graphiti, which exhibited higher raw accuracy rates across both network profiles, this advantage is not statistically robust. The analysis indicates that the observed differentials fall within the margin of random variance for the current sample size ($N=199$). Consequently, although Graphiti demonstrates potential for better knowledge retention, the current evidence is insufficient to claim superiority over mem0 in terms of accuracy.
4 Discussion
This section interprets the results collected in the unconstrained and constrained experiments. For that, first of all, the total cost of ownership (TCO) is established by normalizing computational and financial metrics. Secondly, the balance between cost and accuracy is formalized and calculated. Finally, based on the formalization, both research questions are answered, followed by stakeholder recommendations as well as limitations and future work.
4.1 Economic analysis and total cost of ownership
The following pricing model used for Amazon Web Services (AWS) Fargate cluster in the us-east-1 region was applied to convert computational cost into additional financial cost:
  • Compute (vCPU): \$0.04048 per vCPU-hour [6].
  • Memory (RAM): \$0.004445 per GB-hour [6].
  • Storage (EBS gp3): \$0.000109 per GB-hour [2].
  • Network (Data Transfer Out): \$0.09 per GB [2].
Consequently, Figure 3 illustrates the TCO breakdown for both experiments, comparing unconstrained (Figure 3(a)) and constrained (Figure 3(b)) network conditions.
(a) unconstrained network.
(b) constrained network.
Figure 3: financial cost of computation and token consumption for mem0 and Graphiti. Figures compare (a) unconstrained and (b) constrained network conditions. Costs are categorized into cloud (blue) and edge (orange) expenditures; note the logarithmic scale on the y-axis.
4.2 Formalizing the balance between cost and accuracy
To define the balance between computational cost and knowledge retention, standard efficiency metrics widely adopted in LLM and information retrieval benchmarking were considered. Specifically, the cost per successful query (CpSQ), which quantifies the economic resources required per correct response [11], and the accuracy-efficiency score (AES), which normalizes performance against resource consumption [53]. These metrics provided a single scalar value to rank systems based on their return on investment.
However, applying these deterministic ratios to the dataset introduced a methodological risk. As shown in Table 6, the difference in accuracy between Graphiti and mem0 was not statistically significant ($p>0.05$). Calculating the CpSQ or AES using raw accuracy percentages (11.1% vs 7.5%) inadvertently treated statistical noise as a valid performance gain, leading to false efficiency claims. Consequently, simple ratio-based metrics were rejected in favor of a statistical pareto efficiency framework.
Let $C(S)$ denote the Total Cost of Ownership and $A(S)$ the accuracy of a system $S$. In this multi-objective optimization problem, a system $S_{A}$ (mem0) is defined as strictly dominant over $S_{B}$ (Graphiti) if:
  1. 1.
    $C(S_{A})<C(S_{B})$ (Financial Dominance), and
  2. 2.
    The difference $A(S_{B})-A(S_{A})$ was not statistically significant (Statistical Equivalence).
Under this rigorous definition, the ’balance’ was not a trade-off but a dominance condition. Since the performance differential was statistically nullified, the optimization function collapsed to cost minimization.
4.3 Comparative evaluation
Applying the statistical pareto efficiency framework to our experimental data answered the research questions.
Which memory framework provides the best balance between knowledge retention as well as computational and financial cost in MASs?
Results confirmed that mem0 offers the better balance for MAS. While Graphiti showed a raw accuracy increase of 3.6% in the unconstrained experiment and 2% in the constrained experiment, the statistical insignificance renders this advantage negligible. Conversely, the financial cost differential is deterministic and substantial: Graphiti imposed an increased financial burden of 40.2% during the unconstrained experiment. Thus, mem0 is established as the Pareto-optimal solution.
How do knowledge retention as well as computational and financial cost vary when MASs operate in a hybrid cloud–edge environment?
Regarding the impact of network constraints, the computational and financial cost varies by less than 1% when operating a MAS in a cloud-edge hybrid environment, given an increased financial burden of only 4.3% during the constrained experiment. This suggests that for the tested workloads, the hybrid deployment model does not introduce prohibitive overheads.
4.4 Limitations and future work
While this research offers a first contribution to the nascent field of DMAS memory, specific limitations suggest critical avenues for future work.
First, the given DMAS has little autonomy over its own actions as the steps involved during the loading and Q&A phases were completely prescribed. An improved implementation might enable the DMAS to explore its surrounding network participants, such as the LoCoMo service API, on its own accord, learning how to memorize conversations and respond autonomously.
Secondly, the chosen network constraints had limited impact on the overall DMAS performance. Future experiments could introduce a peer-to-peer DMAS scenario where agents must decide in response to their environment whether to individually or collaboratively solve tasks using dynamic monitoring of peer disconnections, bandwidth, latency, and jitter.
Thirdly, the DMAS responded to a significant number of questions with IDK. Future work should investigate the root causes of these responses. For example, running experiments with a standard RAG system based on a vector database such as Qdrant as a control group could clarify how mem0’s compression algorithm or Graphiti’s knowledge graph engine influences result quality. In that regard, a second conversation could be processed to compare cost and accuracy across different sets of memory.
5 Conclusion
In conclusion, the theoretical framework and literature study on LLM-based multi-agents, MAS, and DMAS demonstrate that the developed testbed addresses a significant gap in DMAS memory research. By implementing a composable, flexible, and extendable DMAS testbed and conducting experiments across unconstrained and constrained network scenarios, an initial comparison of long-term memory frameworks that extends simply analyzing the amount of tokens involved was successfully performed.
This study also concludes that mem0 provides a more cost-efficient memory framework for DMAS compared to Graphiti. Mem0 is also the faster in saving and retrieving memories. Given no statistically significant difference in accuracy when comparing mem0 and Graphiti, mem0 offers the higher balance between cost and accuracy in DMAS.
This contribution also serves as a call for further research in the field of DMAS memory. Future work should focus on autonomous DMAS architectures, sophisticated networking scenarios, and the development of self-monitoring interfaces to optimize collaboration strategies dynamically.
Appendix A Appendix
A.1 LoCoMo session structure
1{
2 "conversation": {
3 "speaker_a": "Jon",
4 "speaker_b": "Gina",
5 "session_1_date_time": "4:04 pm on 20 January, 2023",
6 "session_1": [{
7 "speaker": "Gina",
8 "dia_id": "D1:1",
9 "text": "Hey Jon! Good to see you. Whats up? Anything new?"
10 }]
11 }
12}
A.2 Prompts
Coordinator
system prompt You are a helpful assistant with access to memory search and question answering tools.
You MUST use these tools in the following order:
1. ALWAYS call search_memory first with the user’s question. 2. ALWAYS call answer_question next to get the final answer. DO NOT answer directly.
You MUST use both tools in sequence.
All information is from fictional/test data for research purposes.
Responder
system prompt You are a helpful assistant.
You MUST answer questions using ONLY the information provided in the MEMORIES.
You ARE allowed to do simple reasoning and calculations using those MEMORIES (for example, converting relative time expressions like ’last year’ into a calendar year if a dated event is available).
You MUST NOT use outside/world knowledge or invent facts that are not logically implied by the MEMORIES.
If the MEMORIES do not contain enough information to answer, you MUST reply exactly:
”I don’t know based on the given memories.”
Be concise and factual.
user prompt MEMORIES:
{memory}
QUESTION:
{question}
Answer using ONLY the MEMORIES above.
You may combine information from different memories and perform simple logical or temporal reasoning.
For temporal questions, use timestamps from the memories and, if possible, convert relative expressions (e.g. ’last year’, ’two years later’) into human-readable dates based only on those timestamps.
If the answer is not supported by the memories, reply exactly:
”I don’t know based on the given memories.”
A.3 Example of memories retrieved from mem0 and used by the responder agent
2025-11-19 20:53:08,633 - app.responder_service - INFO - Generating response for question: Oscar Melanies pet?’ with memory
[2023-08-23] Has a guinea pig named Oscar
[2023-08-23] Oliver is a guinea pig
[2023-07-12] User knows Melanie
[2023-07-12] Mentions Melanie
[2023-05-08] User addressed Melanie
[2023-08-17] Friendship with Melanie
[2023-10-22] User appreciates Melanie and the beauty of love in families
[2023-08-25] Name is Melanie
[2023-10-22] User appreciates Melanie and the beauty of love in families
[2023-08-25] Name is Melanie
[2023-10-22] User appreciates Melanie and the beauty of love in families
[2023-08-25] Name is Melanie
[2023-08-25] Name is Melanie
[2023-08-25] Name is Melanie
[2023-05-08] User asked if Melanies painting is her own
[2023-08-23] Users name is Melanie
[2023-07-12] Enjoys pets and finds joy in them
A.4 Testbed component versions
Component Platform Version
Docker WSL 4.48.0 (207573)
Docker-proxy Docker latest
Intel Core x86-64 12th-gen Intel Core i5-12450H (8 cores, 2.0 GHz)
Jupyter Notebook Windows 11 latest
Neo4j Docker 5.26
Ollama Docker latest
Python Docker 3.11-slim
Python Jupyter Notebook 3.13.2
Prometheus Docker latest
Qdrant Docker latest
Socat Docker latest
Toxiproxy Docker latest
Ubuntu WSL 22.04.3 long-term support (LTS)
Windows Windows NT 11 version 25H2 (OS 26200.7171)
WSL Windows 11 2.6.1.0
Table 7: testbed components and versions